OpenStudy (nikosis01):

Help please. Thanks in advance. In March, Isaiah and Freddie had the same amount of money in their savings accounts. In April, Isaiah deposited $210 into his account. Freddie increased the money in his account by 15%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in March

OpenStudy (texaschic101):

m = beginning balance m + 210 = m + 0.15m 210 = 0.15m 210/0.15 = m 1400 = m check.. m + 210 = m + 0.15m 1400 + 210 = 1400 + 0.15(1400) 1610 = 1400 + 210 1610 = 1610 (correct) beginning balance (March) for Isaiah and Freddie was $ 1400.....and in April. it was $ 1610